These are the news starting 01 Jan 2010
Click here for the news of the previous years.

29 Dec 10 Solar irradiance data again on-line.
Starting today at 14:30 solar radiation data from the Vantage Pro Plus backup are displayed and stored at ch. 6 (Solarkip), replacing the readings from the malfunctioning Kip&Zonen pyranometer.
The readings of both sensors are tolerably close, certainly so during this winter period of low solar intensity. For more details, see the last link at the 22Dec10 news.
Relevant graphs will be ok tomorrow ("today") and in 7days ("previous").
23 Dec 10 Solar radiation data problem (continued):
1. On the Live data page (today_01.html) the link to the live Vantage Pro Backup station was corrupt (due to a change of the webserver; my fault!). This has been corrected, and the link "current" under the heading "BKP STATION" shows current Vantage Pro data, among them solar radiation.
2. The batch file making the ddmmyy.txt file had a problem. Seems to be ok now....
22 Dec 10 Attention: Solar pyranometer (ch.6) went bad this morning. Repair probably impossible in the short run.
I programmed a quick kludge to access the data of the Vantage Pro backup station (previous day data):
An example: If today is 22DEC2010, the data from the previous day (which contains total solar irradiance) may be accessed by:
In general:
You are on your own with this!
This short paper might be of interest:
21 Dec  10 Main Linux computer did not restart properly yesterday, sorry for the missing live data. Resumed today at 08:00 UTC. No data lost.
20 Dec 10 1. Maintenance an recalibration work
2. Made backup images from several computers --> no live data from 09:00 to 12:30
08 Nov 10 A couple of weeks ago Eduardo Ferreyra, the president of the FAEC = Argentine Foundation of Ecology visited meteoLCD. Here are his very positive and friendly comments (in Spanish)
01 Nov 10 Uploaded September and October data into the data archive. Sorry for the delay with sep10.dat.
28 Oct 10 The computer managing the radioactivity sensor went bad a couple of days ago. Repaired.
17 Oct 10
I was informed only today that Ernst-Georg Beck which whom I had many exchanges, and who visited meteoLCD some time ago, passed away the 22th Sep. 2010. He was co-author of the paper on "Accurate estimation of CO2..." which we wrote for the Klima2009 conference in Nov. 2009. He will be deeply missed. Two obituaries in German can be found here and here, one in English here.
12 Oct 10 All gas sensors were connected to big UPS (up until now only computers were protected by UPS).. Spikes in data series to temporarly powering down.
27 Sep 10 Maintenance by Envitec. NOx sensor repaired and back on-line.
22 Sep 10 There is a new weblog by F. Massen on moist enthalpy on Prof. Roger Pielke's blog "Climatescience".
Look here for the original contribution and here for a PDF version.
16 Sep 10 1. Thermal regulation of UVB sensor back to normal
2. Moist enthalpy: formula used has been slightly modified (link to full article asap) and y axis has been reduced to [0..80] for better precision.. Near-live graphs now out of test phase.
14 Sep 10 The thermal regulation (Peltier cooler/heater, see channel 18 ) of the UVB sensor behaves erratically since the 06Sep10 18:00. If the problem cannot be solved, the sensor will be replaced.
11 Sep 10 1. I am back from 3 weeks hiking in the Rocky Mountains and doing grand-parent job for a week.
2. The whole equipment ran fine, with the exception of the webcam-proxy computer, who lost power due to loose power cord (hmm...); probably not many did notice this!
3. Maintenance and recalibration work. Upload of September data to data archive asap.
06 Aug 10 Uploaded jul10.dat to data archive.
05 Aug 10 Maintenance and calibration work on the CO2 and O3 sensors.
Rain-gauge was blocked by dirt (since ?? days); cleaned.
22 Jul
I wrote a short article on the measurement of moist enthalpy which is published today as a guest post on Prof. Roger Pielke's blog ClimateScience (see here). To illustrate, I added the graph of moist enthalpy to the page of the near-live data. Read here a comment by Prof. Fucaloro, Claremont Colleges, California.
16 Jul
Prof. Roger Pielke Sr. has added to his well-known blog a short comment of mine concerning moist enthalpy and WBGT measures (see here). I added to the live-data window the graphs (of today and previous 7 days) of the moist enthalpy (in KJ/kg). As calculations are a bit fiddly, consider this for the moment as being in beta test.
09 Jul 10 Calibration and maintenance work on all sensors.
07 Jul 10 Uploaded June 2010 data to data archive.
30 Jun 10 1. There is a serious problem with the NOx sensor since 24Jun10: the internal ozone generator seems to fail. If I am unable to repair, the NOx sensor will be shut down until 17Sep10, date of the next large Envitec maintenance.
I found the fault: the flow rate through the instrument is too low. All the usual controls/replacements/checks do not give a reliable continuous operation, so the AC31M is switched off at 08:50.
2. Added Bezier smoothing curve to dobson graph
18 Jun 10 Maintenance, calibration and repair work by Envicontrol.
12 Jun 10 There was an interruption in the data transfer from the meteoLCD station to the Restena webserver starting the 10th Jun 2010. The cause was a change of name I was unaware of. OK again since today  No data are lost.
06 Jun 10 Uploaded May 2010 data to archive.
10 May 10 Maintenance and calibration work.
02 May 10 Uploaded April 2010 data to data archive. First short but distinctive thunderstorm of the year.
29 Apr 10 WBGT heatstress sensor reactivated today at 08:00 UTC
05 Apr 10 Uploaded March 2010 data to archive.
11 Mar 10 Urgent work at the computers: there was a somewhat obscure NIC problem with the replaced logger computer. Images made from several computers. Big UPS (5000VA) batteries replaced; this is heavy work.
05 Mar 10 1 LAN problem with our DMZ (starting after 07:00).Repaired.
2. Main computer ( = computer making the downloads from the logger) HD died also after 07:00 today (don't ask why...) Replaced computer.
3. UVA sensor is back from recalibration at Solar Light. Reinstalled instrument and adjusted relevant logger factor at 10:00
03 Mar 10 Replaced aging UV lamp of the API ozone sensor (ground ozone) whit a new one (09:30)
02 Mar 10 Some teething problems with the new lightning computer (software and parametrization). OK now.
01 Mar 10 The lightning computer was down since Nov. 2009. Today Jean Mootz and myself installed a new computer (an older refurbished FSC E600) with the latest version of the Lightning2000 software from Aninoquisi..
28 Feb 10 Storm Cynthia caused a power outage; because of missing UPS, an fuse blew when power came back. Repaired at 17:00.
23 Feb 10 The O3 sensor UV lamp problem showed up again, just the day I left for a week abroad. I managed today to solve the problem by carefully repositioning the lamp, which nevertheless is close to agonizing and must be replaced asap.
08 Feb 10 1. Ozone sensor went bad the 5th Feb10 (no output). The cause is probably the UV lamp reaching the lowest permitted limit of luminosity. I checked this morning: actual reading = 2498mV, lowest permitted = 2500mV. After some time, and opening the case to improve cooling, reading went up to 2503mV and output is again available. But this is dangerous living, and the lamp must be replaced asap.
2. UVA sensor is back to Solar Light for recalibration. No UVA data for the moment.
3. There was a power outage at ~13:30; when power came back, a fuse blew and all computers and some sensors came down. Ok again at approx. 16:00.
07 Feb 10 Uploaded jan10.dat to data archive.
15 Jan 10 1. Maintenance work by Envicontrol
2. Still an unsolved Gnuplot problem with the 7 days graphs...
13 Jan 10 1. I repaired the meteo04 Linux computer and put it again online. Be prepared for some disorder in the icons or live graphs...
2. A further problem was a change in the newest 4.2 version of Gnuplot concerning line styles, so all the scripts had to be updated (15:30). The seven days graphs will  (probably) be ok tomorrow.
02 Jan 10 1. The malfunction of several equipments and the communication was due to a big UPS going dead. Not many data are lost (only those whose sensors rely on the UPS outlet). Reetablished working at 09:30.
2. If something goes wrong, more will follow! The hard disk of the main Linux computer is dead-sick. I replaced this computer with a very old one. Live data are running again, but be careful, the scripts are old and I had not the time to recheck everything...
01 Jan 10 1. The usual new year problem with the 7 days graphs.. as the logger file does not contain the year, Gnuplot gets confused the 1rst week of every new year. This will self correct the 7th January.
2. LAN problem started at 03:00 today. Will be repaired asap. No life data for the moment!


Back Home